# Towards Fair RAG: On the Impact of Fair Ranking in Retrieval-Augmented Generation

### v2

@@
Abstract
Many language models now enhance their responses with retrieval capabilities, leading to the widespread adoption of retrieval-augmented generation (RAG) systems. However, despite retrieval being a core component of RAG, much of the research in this area overlooks the extensive body of work on fair ranking, neglecting the importance of considering all stakeholders involved. This paper presents the first systematic evaluation of RAG systems integrated with fair rankings. We focus specifically on measuring the fair exposure of each relevant item across the rankings utilized by RAG systems (i.e., item-side fairness), aiming to promote equitable growth for relevant item providers. To gain a deep understanding of the relationship between item-fairness, ranking quality, and generation quality in the context of RAG, we analyze nine different RAG systems that incorporate fair rankings across seven distinct datasets. Our findings indicate that RAG systems with fair rankings can maintain a high level of generation quality and, in many cases, even outperform traditional RAG systems, despite the general trend of a tradeoff between ensuring fairness and maintaining system-effectiveness. We believe our insights lay the groundwork for responsible and equitable RAG systems and open new avenues for future research. We publicly release our codebase and dataset.  [^1] https://github.com/kimdanny/Fair-RAG
1 Introduction
@@
In recent years, the concept of fair ranking has emerged as a critical concern in modern information access systems [12]. However, despite its significance, fair ranking has yet to be thoroughly examined in the context of retrieval-augmented generation (RAG) [29, 1], a rapidly advancing trend in natural language processing (NLP) systems [27]. To understand why this is important, consider the RAG system in Figure 1, where a user asks a question about running shoes. A classic retrieval system might return several documents containing information from various running shoe companies. If the RAG system only selects the top two documents, then information from the remaining two relevant companies will not be relayed to the predictive model and will likely be omitted from its answer. The fair ranking literature refers to this situation as unfair because some relevant companies (i.e., in documents at position 3 and 4) receive less or no exposure compared to equally relevant company in the top position [12].
@@
Figure 1: Fairness concerns in RAG. A simplified example of how RAG models can ignore equally relevant items ($d_{3}$ and $d_{4}$) and always consume the fixed top-scoring items ($d_{1}$ and $d_{2}$) with the same order of ranking over the multiple user requests. This is due to the deterministic nature of the retrieval process and a short context-length of a language model that necessitates the top-$k$ truncation of a ranked list.
@@
Understanding the effect of fair ranking in RAG is fundamental to ensuring responsible and equitable NLP systems. Since retrieval results in RAG often underlie response attribution [15], unfair exposure of content to the RAG system can result in incomplete evidence in responses (thus compromising recall of potentially relevant information for users) or downstream representational harms (thus creating or reinforcing biases across the set of relevant entities). In situations where content providers are compensated for contributions to inference, there can be financial implications for the unfairness [4, 31, 19]. Indeed, the fair ranking literature indicates that these are precisely the harms that emerge when people are searchers [12], much less RAG systems, where the searchers are machines. RAG complicates these challenges since it often truncates rankings to much shorter lengths to fit the generator’s limited context size [2, 20, 27], making equal exposure of relevant items even harder.
@@
Moreover, the fact that machines are now the searchers necessitates a different notion of item-worthiness (how deserving an item is to be included in a ranked list). Traditionally, ranking quality has been assessed based on relevance labels, which are created according to how relevant an item is to the user’s query [48]. However, with RAG systems, where the consumer is a language model, there is a growing shift towards evaluating ranking quality based on utility labels, which are determined by the usefulness of an item in aiding the model’s task performance, rather than its relevance to the query [43, 57].
@@
This shift from relevance to utility in the concept of item-worthiness can significantly alter our understanding of the relationship between fairness and ranking quality [3]—particularly the tradeoffs that are well-known in the fair ranking literature [5, 10, 51]. Since previous fair ranking studies were conducted based on relevance judgments, they may need to be reexamined in light of utility-based judgments within the context of RAG.
@@
Our research aims to bridge the gap between traditional fair ranking studies and the emerging changes posed by RAG systems, ultimately enhancing our understanding of the interplay between fairness, ranking quality, and the effectiveness of RAG systems. We do this by evaluating RAG systems with a fairness-aware retriever across seven different tasks, experimenting with varying levels of retrieval fairness to observe changes in ranking quality and generation quality (utility). [^2] Throughout this paper, we use ”utility” and ”generation quality” interchangeably to refer to the downstream effectiveness of RAG models, measured by arbitrary string utility metrics.
@@
Our empirical results show that, in the context of machine users, there also exists an overall trend of fairness-quality tradeoff with respect to both retrieval and generation quality. However, the magnitude of this tradeoff is not particularly severe. In fact, we find that RAG models equipped with a fair ranker can often preserve a significant level of retrieval and generation quality, and in some cases, even surpass the quality achieved by the traditional RAG setup with a deterministic ranker that lacks fairness considerations.
@@
2 Background & Related Work
Retrieval-Augmented Generation. RAG, a specific type of retrieval-enhanced machine learning (REML) [54, 27], has been widely adopted in various domains, including language modeling [26], question-answering [22], personalization [46, 45, 28, 35], and recommendation [56]. Studies on the evaluation of RAG models have primarily focused on their effectiveness, including end-to-end performance [22, 18, 29] and the assessment of individual components [14, 42, 43]. However, little research has focused on evaluating fairness in retrieval-enhanced generation models, with the exception of recent work [49], which improved demographic diversity in human image generation by conditioning a generative model with externally retrieved images that help debias the generation process.
Fairness in Ranking. Fair ranking has been approached through various definitions based on normative concerns, primarily with distinctions made according to the stakeholders we prioritize. These include consumer-side fairness [33, 13], which focuses on how fairly a system delivers satisfaction to users; provider-side fairness [47, 23], which addresses how fairly item providers receive monetary or reputational rewards; and item-side fairness [24], which examines how fairly items are treated in terms of representation or exposure. The motivation of item-side fairness is closely linked to provider-side fairness, as unfair treatment of items can lead to unfair compensation for providers. These fairness concerns can be further categorized by the scope of stakeholders, encompassing individual fairness—ensuring similar treatment for similar individuals—and group fairness—ensuring equitable outcomes across different groups [7, 12]. Previous studies have focused on developing metrics to measure fairness [39] and optimizing fair retrievers within a single [52, 55, 47] or multiple rankings [10, 50, 5, 51]. In the context of provider- and item-side fairness, ensuring equal exposure of similar items across multiple rankings has gained significant attention [12]. To achieve this, researchers have used stochastic rankers that return a distribution of rankings, in contrast to deterministic rankers commonly found in areas like RAG, which produce a fixed ranking. This approach ensures that, in expectation, similar items receive equal exposure across multiple user requests, with the distributions typically based on the merit of the rankings, such as an item’s relevance [10, 51].
@@ 2 Background & Related Work
3 Experimental Methodology
Figure 2: Experimental design to investigate the impact of item-fairness on ranking and generation quality in RAG. To evaluate system performance across multiple identical user requests, we sample $N$ rankings from a stochastic retriever. We then measure the fairness and quality of the sampled rankings (Expected Exposure) and assess the system’s expected end-performance (Expected Utility). The query and prompt generators are omitted in the figure for brevity. Details on implementing a RAG system with fair rankings in a production environment can be found in Appendix E.
@@ 3 Experimental Methodology
To address the issue of unfairness in the rankings passed to the generator, we can convert a deterministic retriever into a stochastic retriever, which can, in expectation, provide fair rankings [10]. By sampling a ranking based on its quality to users—in this case generators—the expected exposure of different relevant items becomes similar and, therefore, fairer (Appendix E). Because decisions are stochastic, the fairness and quality of stochastic retrieval is evaluated based on a sample of rankings. Similarly, since a sampled ranking is processed by a generator, we also compute the expected generator effectiveness over sampled rankings. The complete evaluation pipeline of a RAG system with a stochastic retriever is illustrated in Figure 2.
The following sections describe how we construct a test collection with utility labels (§3.1), how we stochastically sample multiple rankings (§3.2), and how we evaluate the fairness and ranking quality of the sampled rankings (§3.3.1) and measure the effectiveness of a RAG system given multiple rankings (§3.3.2).
3.1 Construction of a Test Collection with Utility Labels
Setting an appropriate proxy for measuring item-worthiness is crucial in the evaluation of fairness [3]. Drawing on the insight that utility-based judgments are more suitable than relevance judgments in the context of RAG [57, 43], we annotate item-level utility labels for all items in the corpus.
@@ 3.1 Construction of a Test Collection with Utility Labels
Therefore, the item-level utility labels are designed to be both task- and generator-dependent, as the utility of each item varies depending on the task and the language model used. This labeling process also aligns with the principles of task-based information retrieval, where, in the context of human searchers, document utility may vary on how the user expects to use the document [25].
3.2 Fairness-Aware Stochastic Retriever
Stochastic retrievers have been used for various purposes, such as optimization of retrieval models [6, 53, 16, 36], as well as ensuring equitable exposure of items [37, 10, 36]. Many of these studies use Plackett-Luce sampling [38] to achieve the stochasticity of retrieval. We follow the line of research and formally define how we derive a fairness-aware stochastic retriever through Plackett-Luce sampling. To enhance sampling efficiency, we adopt the methodology of Oosterhuis [36], and for controllable randomization, we utilize the approach proposed by Diaz et al. [10].
Given $n$ items in a corpus C, a vector of retrieval scores $\mathbf{s}\in\mathbb{R}^{n}$ can be obtained from $\mathcal{R}(q,\textrm{C})$, which can be used to generate a ranked list $L$. We then min-max normalize retrieval scores to be in [0, 1] in order to construct a multinomial distribution over items [5]. The probability of an item $d$ being selected as the $i$’th item in a new ranking $\pi$ through Plackett-Luce sampling is given by
$p(d|L_{1:i-1})=\frac{\text{exp}(\bar{\mathbf{s}}_{d})\mathds{1}[d\notin L_{1:i-1}]}{\sum_{d^{\prime}\in\textrm{C}\backslash L_{1:i-1}}\text{exp}(\bar{\mathbf{s}}_{d^{\prime}})}$ (1)
@@ 3.2 Fairness-Aware Stochastic Retriever
This repeated sampling and renormalization process can be efficiently managed using the Gumbel-Softmax trick [17, 32], which enables the sampling of rankings to be performed at the speed of sorting [36]. To do so, for each sampling iteration, we draw $U_{i}\sim\text{Uniform}(0,1)$, followed by generating a Gumbel noise $G_{i}=-\log(-\log(U_{i}))$. The probability of each sampled ranking is then obtained by sorting the items based on their perturbed scores $\tilde{\mathbf{s}}_{d_{i}}=\bar{\mathbf{s}}_{d_{i}}+G_{i}$.
3.2.1 Controlling the Level of Fairness
Adjusting the level of randomization directly controls the degree of item-fairness, aligning with our goal to observe how varying levels of fairness in rankings affect the ranking and generation quality of a RAG model. To obtain the controllability, we follow the work of Diaz et al. [10] and use a fairness control parameter $\alpha$. We apply the scalar $\alpha$ to each value in the normalized score vector $\bar{\mathbf{s}}$ by raising each value to the power of $\alpha$ [^3] We normalized the values to the range of [1, 2] instead of [0, 1]. The addition of 1 effectively serves the same purpose as adjusting a real-numbered $\alpha$. We chose this range to allow for an integer-valued $\alpha$. This process is done before the scores are passed to the sampling policy. Therefore, the modified sampling distribution is thus defined as:
$p(d|L_{1:i-1})=\frac{\text{exp}(\bar{\mathbf{s}}^{\alpha}_{d})\mathds{1}[d\notin L_{1:i-1}]}{\sum_{d^{\prime}\in\textrm{C}\backslash L_{1:i-1}}\text{exp}(\bar{\mathbf{s}}^{\alpha}_{d^{\prime}})}$ (2)

### v3

@@
Abstract.
Modern language models frequently include retrieval components to improve their outputs, giving rise to a growing number of retrieval-augmented generation (RAG) systems. Yet, most existing work in RAG has underemphasized fair ranking techniques and neglected the diverse interests of all stakeholders. In this paper, we present the first comprehensive study of RAG systems that incorporate fairness-aware rankings, focusing on both ranking fairness and attribution fairness—ensuring equitable exposure of sources cited in the final text. We specifically examine item-side fairness, i.e., whether retrieved documents receive balanced exposure, and assess how this affects both the system’s overall performance and the eventual distribution of cited sources. Across twelve RAG models and seven tasks, we find that fairness-aware retrieval frequently retains or even improves ranking effectiveness and generation quality, countering the widespread belief that fairness compromises system performance. Moreover, we show that fair retrieval leads to more balanced attribution in the final responses, ensuring that the cited sources are credited more equitably. Our results underscore the importance of item-side fairness throughout both retrieval and generation phases, offering key insights for building more responsible and equitable RAG systems and illustrating promising avenues for future exploration in fair ranking and source attribution.  [^1] We publicly release our codebase and dataset at https://github.com/kimdanny/Fair-RAG.
Fair Ranking, Fair Attribution, Retrieval-Augmented Generation
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2025; Woodstock, NYisbn: 978-1-4503-XXXX-X/2025/06ccs: Computing methodologies Natural language generationccs: Information systems Evaluation of retrieval results
1. Introduction
@@
In recent years, the concept of fair ranking has emerged as a critical concern in modern information access systems (Ekstrand et al., 2022). However, despite its significance, fair ranking has yet to be thoroughly examined in the context of retrieval-augmented generation (RAG) (Lewis et al., 2020; Asai et al., 2024), a rapidly advancing trend in natural language processing (NLP) systems (Kim et al., 2024). To understand why this is important, consider the RAG system in Figure 1, where a user asks a question about running shoes. A classic retrieval system might return several documents containing information from various running shoe companies. If the RAG system only selects the top two documents, then information from the remaining two relevant companies will not be relayed to the predictive model and will likely be omitted from its answer. The fair ranking literature refers to this situation as unfair because some relevant companies (i.e., in documents at position 3 and 4) receive less or no exposure compared to equally relevant company in the top position (Ekstrand et al., 2022).
@@
Figure 1. Fairness concerns in RAG. A simplified example of how RAG models can ignore equally relevant items ($d_{3}$ and $d_{4}$) and always consume the fixed top-scoring items ($d_{1}$ and $d_{2}$) with the same order of ranking over the multiple user requests. This is due to the deterministic nature of the retrieval process and a short context-length of a language model that necessitates the top-$k$ truncation of a ranked list.
@@
Understanding the effect of fair ranking in RAG is fundamental to ensuring responsible and equitable NLP systems. Since retrieval results in RAG often underlie response attribution (Gao et al., 2023b), unfair exposure of content to the RAG system can result in incomplete evidence in responses (thus compromising recall of potentially relevant information for users) or downstream representational harms (thus creating or reinforcing biases across the set of relevant entities). In situations where content providers are compensated for contributions to inference, there can be financial implications for the unfairness (Balan et al., 2023; Lyu et al., 2023; Henderson et al., 2023). Indeed, the fair ranking literature indicates that these are precisely the harms that emerge when people are searchers (Ekstrand et al., 2022), much less RAG systems, where the searchers are machines. RAG complicates these challenges since it often truncates rankings to much shorter lengths to fit the generator’s limited context size (Bahri et al., 2020; Hofstätter et al., 2023; Kim et al., 2024), making equal exposure of relevant items even harder.
@@
Moreover, the fact that machines are now the searchers necessitates a different notion of item-worthiness (how deserving an item is to be included in a ranked list). Traditionally, ranking quality has been assessed based on relevance labels, which are created according to how relevant an item is to the user’s query (Saracevic, 2016). However, with RAG systems, where the consumer is a language model, there is a growing shift towards evaluating ranking quality based on utility labels, which are determined by the usefulness of an item in aiding the model’s task performance, rather than its relevance to the query (Salemi and Zamani, 2024; Zhang et al., 2024b).
@@
This shift from relevance to utility in the concept of item-worthiness can significantly alter our understanding of the relationship between fairness and ranking quality (Balagopalan et al., 2023)—particularly the tradeoffs that are well-known in the fair ranking literature (Biega et al., 2018; Diaz et al., 2020; Singh and Joachims, 2019). Since previous fair ranking studies were conducted based on relevance judgments, they may need to be reexamined in light of utility-based judgments within the context of RAG.
However, purely focusing on how often certain items appear in the top-$k$ positions can neglect the fact that not all retrieved items are necessarily attributed in the final generated response. If an item is retrieved but never actually influences the RAG model’s output, one cannot fully gauge whether it truly received exposure from the standpoint of the final generation. This reveals a subtle yet important gap: fair retrieval may not directly translate to fair consumption, especially when some retrieved items might be overshadowed in the generation step. Measuring how exposure is ultimately distributed across the attributed sources in the final response offers a more complete picture of exposure-based fairness in the context of RAG systems.
@@
Our research aims to bridge the gap between traditional fair ranking studies and the emerging changes posed by RAG systems, ultimately enhancing our understanding of the interplay between fairness, ranking quality, and the effectiveness of RAG systems. We do this by evaluating RAG systems with a fairness-aware retriever across seven different tasks, experimenting with varying levels of retrieval fairness to observe changes in ranking quality and generation quality (utility) [^2] Throughout this paper, we use ”utility” and ”generation quality” interchangeably to refer to the downstream effectiveness of RAG models, measured by arbitrary string utility metrics., as well as the fairness of attributed sources.
@@
Our empirical results show that, in the context of machine users, there also exists an overall trend of fairness-quality tradeoff with respect to both retrieval and generation quality. However, the magnitude of this tradeoff is not particularly severe. In fact, we find that RAG models equipped with a fair ranker can often preserve a significant level of retrieval and generation quality, and in some cases, even surpass the quality achieved by the traditional RAG setup with a deterministic ranker that lacks fairness considerations. Moreover, while the fraction of retrieved sources that actually appear in the final response may vary, equitable retrieval frequently leads to more equitable usage of those sources by the generator.
@@
2. Background & Related Work
2.1. Retrieval-Augmented Generation
RAG, a specific type of retrieval-enhanced machine learning (REML) (Zamani et al., 2022; Kim et al., 2024), has been widely adopted in various domains, including language modeling (Khandelwal et al., 2020), question-answering (Izacard et al., 2023), and personalization (Salemi et al., 2024b, a; Neelakanteswara et al., 2024). Studies on the evaluation of RAG models have primarily focused on their effectiveness, including end-to-end performance (Izacard et al., 2023; Guu et al., 2020; Lewis et al., 2020) and the assessment of individual components (Es et al., 2024; Saad-Falcon et al., 2024; Salemi and Zamani, 2024; Ru et al., 2024), such as retrieval relevance and model faithfulness. Furthermore, recent efforts have explored attribution mechanisms for ensuring the trustworthiness of RAG responses (Gao et al., 2023a, b), examining how thoroughly models reference the source text and thus promoting a more faithful generation process.
However, little research has focused on evaluating fairness in retrieval-enhanced generation models, with the exception of recent work (Shrestha et al., 2024), which improved demographic diversity in human image generation by conditioning a generative model with externally retrieved images that help debias the generation process.
2.2. Fairness in Ranking
Fair ranking has been approached through various definitions based on normative concerns, primarily with distinctions made according to the stakeholders we prioritize. These include consumer-side fairness (Mehrotra et al., 2017; Ekstrand et al., 2024), which focuses on how fairly a system delivers satisfaction to users; provider-side fairness (Sapiezynski et al., 2019; Jaenich et al., 2024), which addresses how fairly item providers receive monetary or reputational rewards; and item-side fairness (Jiang et al., 2024), which examines how fairly items are treated in terms of representation or exposure. The motivation of item-side fairness is closely linked to provider-side fairness, as unfair treatment of items can lead to unfair compensation for providers. These fairness concerns can be further categorized by the scope of stakeholders, encompassing individual fairness—ensuring similar treatment for similar individuals—and group fairness—ensuring equitable outcomes across different groups (Caton and Haas, 2020; Ekstrand et al., 2022). Previous studies have focused on developing metrics to measure fairness (Raj and Ekstrand, 2020) and optimizing fair retrievers within a single (Yang and Stoyanovich, 2017; Zehlike et al., 2017; Sapiezynski et al., 2019) or multiple rankings (Diaz et al., 2020; Singh and Joachims, 2018; Biega et al., 2018; Singh and Joachims, 2019). In the context of provider- and item-side fairness, ensuring equal exposure of similar items across multiple rankings has gained significant attention (Ekstrand et al., 2022). To achieve this, researchers have used stochastic rankers that return a distribution of rankings, in contrast to deterministic rankers commonly found in areas like RAG, which produce a fixed ranking. This approach ensures that, in expectation, similar items receive equal exposure across multiple user requests, with the distributions typically based on the merit of the rankings, such as an item’s relevance (Diaz et al., 2020; Singh and Joachims, 2019).
@@ 2.2. Fairness in Ranking
3. Experimental Methodology
Figure 2. Overview of our experimental design for examining how item-fairness in retrieval impacts both ranking and generation quality, as well as the fairness of attributed items in retrieval-augmented generation. To evaluate system performance across multiple identical user requests, we sample $N$ rankings from a stochastic retriever. Then, we measure the fairness and quality of these rankings (Expected Exposure §3.3.1), and then assess the system’s overall performance (Expected Utility §3.3.2), the usage of retrieved evidence (Expected Attribution Rate §3.3.3), and the fairness of attributed sources (Expected Attributed Exposure §3.3.4). The query and prompt generators are omitted for brevity.
@@ 3. Experimental Methodology
To address the issue of unfairness in the rankings passed to the generator, we can convert a deterministic retriever into a stochastic retriever, which can, in expectation, provide fair rankings (Diaz et al., 2020). By sampling a ranking based on its quality to users—in this case generators—the expected exposure of different relevant items becomes similar and, therefore, fairer. Because decisions are stochastic, the fairness and quality of stochastic retrieval is evaluated based on a sample of rankings. Likewise, since each sampled ranking is processed by a generator, we measure the expected generator effectiveness over these sampled rankings. We also consider how fairly exposure is allocated among the items that end up being attributed by the final output, acknowledging that some retrieved items may not be credited in the generated text. The complete evaluation pipeline of a RAG system with a stochastic retriever is illustrated in Figure 2.
The following sections describe how we construct a test collection with utility labels (§3.1), how we stochastically sample multiple rankings (§3.2), how we evaluate the fairness and ranking quality of the sampled rankings (§3.3.1), and how we assess a RAG system’s performance over sampled rankings (§3.3.2). We then explain how we measure the expected number of attributed items (§3.3.3) and evaluate the fairness of these attributed sources (§3.3.4).
3.1. Construction of a Test Collection with Utility Labels
Setting an appropriate proxy for measuring item-worthiness is crucial in the evaluation of fairness (Balagopalan et al., 2023). Drawing on the insight that utility-based judgments are more suitable than relevance judgments in the context of RAG (Zhang et al., 2024b; Salemi and Zamani, 2024), we annotate item-level utility labels for all items in the corpus.
@@ 3.1. Construction of a Test Collection with Utility Labels
Therefore, the item-level utility labels are designed to be both task- and generator-dependent, as the utility of each item varies depending on the task and the language model used. This labeling process also aligns with the principles of task-based information retrieval, where, in the context of human searchers, document utility may vary on how the user expects to use the document (Kelly et al., 2013).
3.2. Fairness-Aware Stochastic Retriever
Stochastic retrievers have been used for various purposes, such as optimization of retrieval models (Bruch et al., 2020; Zamani and Bendersky, 2024; Guiver and Snelson, 2009; Oosterhuis, 2021), as well as ensuring equitable exposure of items (Oosterhuis, 2022; Diaz et al., 2020; Oosterhuis, 2021). Many of these studies use Plackett-Luce sampling (Plackett, 1975) to achieve the stochasticity of retrieval. We follow the line of research and formally define how we derive a fairness-aware stochastic retriever through Plackett-Luce sampling. To enhance sampling efficiency, we adopt the methodology of Oosterhuis (2021), and for controllable randomization, we utilize the approach proposed by Diaz et al. (2020).
Given $n$ items in a corpus C, a vector of retrieval scores $\mathbf{s}\in\mathbb{R}^{n}$ can be obtained from $\mathcal{R}(q,\textrm{C})$, which can be used to generate a ranked list $L$. We then min-max normalize retrieval scores to be in [0, 1] in order to construct a multinomial distribution over items (Biega et al., 2018). The probability of an item $d$ being selected as the $i$’th item in a new ranking $\pi$ through Plackett-Luce sampling is given by
(1) $p(d|L_{1:i-1})=\frac{\text{exp}(\bar{\mathbf{s}}_{d})\mathds{1}[d\notin L_{1:i-1}]}{\sum_{d^{\prime}\in\textrm{C}\backslash L_{1:i-1}}\text{exp}(\bar{\mathbf{s}}_{d^{\prime}})}$
@@ 3.2. Fairness-Aware Stochastic Retriever
This repeated sampling and renormalization process can be efficiently managed using the Gumbel-Softmax trick (Gumbel, 1954; Maddison et al., 2017), which enables the sampling of rankings to be performed at the speed of sorting (Oosterhuis, 2021). To do so, for each sampling iteration, we draw $U_{i}\sim\text{Uniform}(0,1)$, followed by generating a Gumbel noise $G_{i}=-\log(-\log(U_{i}))$. The probability of each sampled ranking is then obtained by sorting the items based on their perturbed scores $\tilde{\mathbf{s}}_{d_{i}}=\bar{\mathbf{s}}_{d_{i}}+G_{i}$.
3.2.1. Controlling the Level of Fairness
Adjusting the level of randomization directly controls the degree of item-fairness, aligning with our goal to observe how varying levels of fairness in rankings affect the ranking and generation quality of a RAG model. To obtain the controllability, we follow the work of Diaz et al. (2020) and use a fairness control parameter $\alpha$. We apply the scalar $\alpha$ to each value in the normalized score vector $\bar{\mathbf{s}}$ by raising each value to the power of $\alpha$ [^3] We normalized the values to the range of [1, 2] instead of [0, 1]. The addition of 1 effectively serves the same purpose as adjusting a real-numbered $\alpha$. We chose this range to allow for an integer-valued $\alpha$. This process is done before the scores are passed to the sampling policy. Therefore, the modified sampling distribution is thus defined as:
(2) $p(d|L_{1:i-1})=\frac{\text{exp}(\bar{\mathbf{s}}^{\alpha}_{d})\mathds{1}[d\notin L_{1:i-1}]}{\sum_{d^{\prime}\in\textrm{C}\backslash L_{1:i-1}}\text{exp}(\bar{\mathbf{s}}^{\alpha}_{d^{\prime}})}$